# The Poisson Midpoint Method for Langevin Dynamics: Provably Efficient Discretization for Diffusion Models

### v2

@@ Appendix D Proof of Lemma 7
$=\sqrt{1+\nu}\mathcal{W}_{2}\left(\langle\mathbf{Z},e_{1}\rangle,\frac{\langle\mathbf{Z}^{\prime},e_{1}\rangle+\langle\mathbf{N},e_{1}\rangle}{\sqrt{1+\nu}}\right)$ (27)
@@
    $|Q_{1}|\leq\frac{m|N_{1}N_{1}^{\prime}|}{m+1}+\frac{\beta^{2}}{m+1}+\frac{1}{2(m^{2}-1)}$ almost surely
@@
    $\mathbb{E}[Q_{1}^{2}]\leq\frac{m\beta^{2}+2m^{2}+1}{2(m^{2}-1)^{2}}$
@@
    $|Q_{1}|$ $\leq\frac{m|N_{1}N_{1}^{\prime}|}{m+1}+\frac{\beta^{2}}{m+1}+\frac{1}{2(m^{2}-1)}$ (29)
@@
$\left[\mathcal{W}_{2}\left(\langle\mathbf{Z},e_{1}\rangle,\frac{\langle\mathbf{Z}^{\prime},e_{1}\rangle+\langle\mathbf{N},e_{1}\rangle}{\sqrt{1+\nu}}\right)\right]^{2}\leq 2[\mathbb{E}e^{Q_{1}}-1]=2\left[\mathbb{E}[Q_{1}]+\frac{1}{2}\mathbb{E}[Q_{1}^{2}]+\mathbb{E}[R(Q_{1})]\right]$ (31)
@@
$\mathbb{E}[Q_{1}]+\frac{1}{2}\mathbb{E}[Q_{1}^{2}]\leq\frac{3+m\beta^{2}}{4(m^{2}-1)^{2}}$
@@
$\leq\left[\frac{m\beta^{4}+m^{2}\beta^{2}}{4(m^{2}-1)^{2}}\right]\exp(\tfrac{3\beta^{2}}{2})$ (33)
@@
$\left[\mathcal{W}_{2}\left(\langle\mathbf{Z},e_{1}\rangle,\frac{\langle\mathbf{Z}^{\prime},e_{1}\rangle+\langle\mathbf{N},e_{1}\rangle}{\sqrt{1+\nu}}\right)\right]^{2}\leq\left[\frac{m\beta^{4}+m^{2}\beta^{2}}{2(m^{2}-1)^{2}}\right]\exp(\tfrac{3\beta^{2}}{2})+\frac{3+m\beta^{2}}{2(m^{2}-1)^{2}}$
@@
$\left[\mathcal{W}_{2}\left(\langle\mathbf{Z},e_{1}\rangle,\frac{\langle\mathbf{Z}^{\prime},e_{1}\rangle+\langle\mathbf{N},e_{1}\rangle}{\sqrt{1+\nu}}\right)\right]^{2}\leq\left[\frac{\beta^{4}\nu^{3}}{2}+\frac{\beta^{2}\nu^{2}}{2}\right]\exp(\tfrac{3\beta^{2}}{2})+\frac{3\nu^{4}+\beta^{2}\nu^{3}}{2}$ (34)
@@ Appendix E Overdamped Langevin Dynamics
Recall $B_{tK+i}$ from Theorem 1. Instantiating this for Overdamped Langevin Dynamics, we conclude that for any $t,i\in\mathbb{N}\cup\{0\}$ and $10\leq i\leq K-$ 1, we must have:
$\|B_{tK+i}\|^{2}\leq\frac{L^{2}\alpha}{2K}\sup_{0\leq j\leq K-1}\|\hat{\tilde{X}}_{tK+j}-\tilde{X}_{tK+j}\|^{2}$ (35)
@@
Now, consider $\beta_{tK+i}$ in Theorem 1. For any $p\geq 1$, $t,i\in\mathbb{N}\cup\{0\}$ and $10\leq i\leq K-$ 1, we have:
$\beta_{tK+i}^{2p}\leq(\tfrac{L^{2}\alpha K}{2})^{p}\sup_{0\leq j\leq K-1}\|\hat{\tilde{X}}_{tK+j}-\tilde{X}_{tK}\|^{2p}$ (36)
@@ E.1 Bounding the Moments
Suppose the stepsize $\alpha L<1$. Let $M_{t,k}:=\sqrt{\frac{2\alpha}{K}}\sup_{0\leq i\leq k-1}\bigr{\|}\sum_{j=0}^{i}Z_{tK+j}\bigr{\|}$. Then the following hold almost surely for every $0\leq k\leq K$
@@
    $\sup_{0\leq i\leq k-1}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}\|\leq\alpha\|\nabla F(\tilde{X}_{tK})\|+M_{t,k}$
@@
    $\sup_{0\leq i\leq k}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK+i}\|\leq\alpha LN_{t}\sup_{i\leq k-1}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}\|$
@@
    $\mathbb{E}[M_{t,K}^{p}]\leq C(p)(\alpha d)^{\frac{p}{2}}$
@@
$\sup_{t\leq h\leq s}\left[\mathbb{E}\|\tilde{X}_{hK}-\tilde{X}_{tK}\|^{p}|\tilde{X}_{tK}\right]^{\tfrac{1}{p}}$ $\leq 3\alpha(s-t)\|\nabla F(\tilde{X}_{tK})\|+C_{p}\sqrt{\alpha d(s-t)}$
@@
$\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}\leq C_{p}L^{p}d^{p}T+C_{p}(\alpha L)^{p-1}\mathbb{E}(\sum_{t=1}^{T}\|\nabla F(\tilde{X}_{tK})\|^{2})^{p}$ (37)
@@
$\quad+C_{p}L^{p}d^{p}T^{p}+\frac{C_{p}}{\alpha^{p}}$ (38)
@@
$\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}\leq C_{p}L^{p}d^{p}T(1+(\alpha LT)^{p-1})+\frac{C_{p}L^{p-1}}{\alpha}\left[\mathbb{E}|(F(X_{0})-F(\tilde{X}_{KT}))^{+}|^{p}+1\right]$
@@
    $\mathbb{E}\|B_{tK+i}\|^{2}\leq\frac{CL^{4}\alpha^{5}}{K}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\frac{CL^{4}\alpha^{4}}{K}d$
@@
    $\mathbb{E}\beta_{tK+i}^{2p}\leq C_{p}\left[L^{2p}\alpha^{3p}K^{p}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}+L^{2p}\alpha^{2p}K^{p}d^{p}\right]$
@@
    $\leq\frac{CL^{4}\alpha^{3}}{K}\left[\alpha^{2}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\alpha d\right]$ (39)
@@ E.2 Finishing The Proof
$\quad+L^{10}\alpha^{10}K^{3}d^{5}T+L^{19}\alpha^{19}K^{3}d^{5}T^{5}+L^{14}\alpha^{14}K^{7}d^{7}T+L^{27}\alpha^{27}K^{7}d^{7}T^{7}$ (40)
@@ Appendix F Underdamped Langevin Dynamics
$A_{h}:=\begin{bmatrix}\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}\\ 0&e^{-\gamma h}\mathbf{I}_{d}\end{bmatrix},\quad G_{h}:=\begin{bmatrix}\tfrac{1}{\gamma}(h-\tfrac{1}{\gamma}(1-e^{-\gamma h}))\mathbf{I}_{d}&0\\ \tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}&0\end{bmatrix}\quad b(X_{t}):=\begin{bmatrix}-\nabla F(U_{t})\\ 0\end{bmatrix}$
@@
$G_{h}^{\intercal}\Gamma_{h}^{-2}G_{h}\preceq\begin{bmatrix}C\frac{h\exp(2\gamma h)}{\gamma}\mathbf{I}_{d}&0\\ 0&0\end{bmatrix}$
For all $10\leq j\leq K-$ 1,
@@ F.1 Bounding the Moments:
Suppose that $\|\cdot\|_{\mathsf{any}}$ is any semi-norm over $\mathbb{R}^{d}$. Then the following hold almost surely for every $0\leq k\leq K$
@@
    $\quad+\|\sum_{j=0}^{i-1}A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}\|_{\mathsf{any}}$
@@
Consider the seminorm given by $\|x\|^{2}_{\Pi}=x^{\intercal}\Pi x$. Let $M_{t,K}:=\sup_{j\leq K-1}\|\sum_{j=0}^{i-1}A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}\|_{\Pi}$. For any $10\leq i\leq K-$ 1.
@@
    $\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|^{2}\leq C\alpha^{2}\|\tilde{V}_{tK}\|^{2}+C\alpha^{4}\|\nabla F(\tilde{U}_{tK})\|^{2}+CM_{t,K}^{2}$ (42)
@@
    $\mathbb{E}M_{t,K}^{p}\leq C\exp(\frac{p\gamma\alpha}{2})\gamma^{\tfrac{p}{2}}\alpha^{\tfrac{3p}{2}}(d+\log K)^{\frac{p}{2}}$
@@
    $\quad+\frac{L^{2}T^{1-\tfrac{1}{p}}\alpha^{2}}{\gamma^{2}}[\mathbb{E}\mathcal{S}_{2p}(V)]^{\tfrac{1}{p}}+\frac{L^{2}T^{1-\tfrac{1}{p}}\alpha^{4}}{\gamma^{2}}[\mathbb{E}\mathcal{S}_{2p}(\nabla F)]^{\tfrac{1}{p}}$ (44)
@@
    $\quad+\frac{L^{2}T^{1-\tfrac{1}{p}}\alpha^{4}}{\gamma^{2}}[\mathbb{E}\mathcal{S}_{2p}(\nabla F)]^{\tfrac{1}{p}}$ (46)
@@
$\quad+C_{p}\sqrt{\frac{d+\log K}{L}}$ (48)
@@
$\quad+C_{p}\gamma\sqrt{\frac{d+\log K}{L}}$ (49)
@@
    $\mathbb{E}\mathcal{S}_{2p}(\nabla F)\leq C_{p}(L\alpha N)^{2p}\mathbb{E}\mathcal{S}_{2p}(V)+C_{p}TL^{p}\left(d+\log K\right)^{p}+\frac{2^{p}}{N^{p-1}}\mathbb{E}(\mathcal{S}_{2}(\nabla F))^{p}$ (50)
@@
    $\mathbb{E}\mathcal{S}_{2p}(V)\leq C_{p}(\alpha N)^{2p}\mathbb{E}\mathcal{S}_{2p}(\nabla F)+T\left[\frac{\gamma^{2}}{L}(d+\log K)\right]^{p}+\frac{2^{p}}{N^{p-1}}\mathbb{E}(\mathcal{S}_{2}(V))^{p}$ (51)
@@
$\quad+C_{p}NTL^{p}\left(d+\log K\right)^{p}$ (52)
@@
$\mathbb{E}\mathcal{S}_{2p}(\nabla F)\leq C_{p}(L\alpha N)^{2p}\mathbb{E}\mathcal{S}_{2p}(V)+C_{p}TL^{p}\left(d+\log K\right)^{p}+\frac{2^{p}}{N^{p-1}}\mathbb{E}(\mathcal{S}_{2}(\nabla F))^{p}$ (53)
@@
$\quad+C_{p}NT\frac{\gamma^{2p}}{L^{p}}\left(d+\log K\right)^{p}$ (54)
@@
$\mathbb{E}\mathcal{S}_{2p}(V)\leq C_{p}(\alpha N)^{2p}\mathbb{E}\mathcal{S}_{2p}(\nabla F)+T\left[\frac{\gamma^{2}}{L}(d+\log K)\right]^{p}+\frac{2^{p}}{N^{p-1}}\mathbb{E}(\mathcal{S}_{2}(V))^{p}$ (55)
@@
$\quad C_{p}T\left[\tfrac{\gamma^{4p}}{L^{p}}+(\gamma\alpha T)^{p-1}\gamma^{2p}\right](d+\log K)^{p}$
@@
$\quad+C_{p}T\left[\tfrac{\gamma^{2p}}{L^{p}}+(\gamma\alpha T)^{p-1}\right](d+\log K)^{p}$

### v3

@@ Appendix D Proof of Lemma 7
$=\sqrt{1+\nu}\mathcal{W}_{2}\left(\langle\mathbf{Z},e_{1}\rangle,\frac{\langle\mathbf{Z}^{\prime},e_{1}\rangle+\langle\mathbf{N},e_{1}\rangle}{\sqrt{1+\nu}}\right).$ (27)
@@
    $|Q_{1}|\leq\frac{m|N_{1}N_{1}^{\prime}|}{m+1}+\frac{\beta^{2}}{m+1}+\frac{1}{2(m^{2}-1)}$ almost surely.
@@
    $\mathbb{E}[Q_{1}^{2}]\leq\frac{m\beta^{2}+2m^{2}+1}{2(m^{2}-1)^{2}}$.
@@
    $|Q_{1}|$ $\leq\frac{m|N_{1}N_{1}^{\prime}|}{m+1}+\frac{\beta^{2}}{m+1}+\frac{1}{2(m^{2}-1)}.$ (29)
@@
$\left[\mathcal{W}_{2}\left(\langle\mathbf{Z},e_{1}\rangle,\frac{\langle\mathbf{Z}^{\prime},e_{1}\rangle+\langle\mathbf{N},e_{1}\rangle}{\sqrt{1+\nu}}\right)\right]^{2}\leq 2[\mathbb{E}e^{Q_{1}}-1]=2\left[\mathbb{E}[Q_{1}]+\frac{1}{2}\mathbb{E}[Q_{1}^{2}]+\mathbb{E}[R(Q_{1})]\right].$ (31)
@@
$\mathbb{E}[Q_{1}]+\frac{1}{2}\mathbb{E}[Q_{1}^{2}]\leq\frac{3+m\beta^{2}}{4(m^{2}-1)^{2}}.$
@@
$\leq\left[\frac{m\beta^{4}+m^{2}\beta^{2}}{4(m^{2}-1)^{2}}\right]\exp(\tfrac{3\beta^{2}}{2}).$ (33)
@@
$\left[\mathcal{W}_{2}\left(\langle\mathbf{Z},e_{1}\rangle,\frac{\langle\mathbf{Z}^{\prime},e_{1}\rangle+\langle\mathbf{N},e_{1}\rangle}{\sqrt{1+\nu}}\right)\right]^{2}\leq\left[\frac{m\beta^{4}+m^{2}\beta^{2}}{2(m^{2}-1)^{2}}\right]\exp(\tfrac{3\beta^{2}}{2})+\frac{3+m\beta^{2}}{2(m^{2}-1)^{2}}.$
@@
$\left[\mathcal{W}_{2}\left(\langle\mathbf{Z},e_{1}\rangle,\frac{\langle\mathbf{Z}^{\prime},e_{1}\rangle+\langle\mathbf{N},e_{1}\rangle}{\sqrt{1+\nu}}\right)\right]^{2}\leq\left[\frac{\beta^{4}\nu^{3}}{2}+\frac{\beta^{2}\nu^{2}}{2}\right]\exp(\tfrac{3\beta^{2}}{2})+\frac{3\nu^{4}+\beta^{2}\nu^{3}}{2}.$ (34)
@@ Appendix E Overdamped Langevin Dynamics
Recall $B_{tK+i}$ from Theorem 1. Instantiating this for Overdamped Langevin Dynamics, we conclude that for any $t,i\in\mathbb{N}\cup\{0\}$ and $0\leq i\leq K-1$, we must have:
$\|B_{tK+i}\|^{2}\leq\frac{L^{2}\alpha}{2K}\sup_{0\leq j\leq K-1}\|\hat{\tilde{X}}_{tK+j}-\tilde{X}_{tK+j}\|^{2}.$ (35)
@@
Now, consider $\beta_{tK+i}$ in Theorem 1. For any $p\geq 1$, $t,i\in\mathbb{N}\cup\{0\}$ and $0\leq i\leq K-1$, we have:
$\beta_{tK+i}^{2p}\leq(\tfrac{L^{2}\alpha K}{2})^{p}\sup_{0\leq j\leq K-1}\|\hat{\tilde{X}}_{tK+j}-\tilde{X}_{tK}\|^{2p}.$ (36)
@@ E.1 Bounding the Moments
Suppose the stepsize $\alpha L<1$. Let $M_{t,k}:=\sqrt{\frac{2\alpha}{K}}\sup_{0\leq i\leq k-1}\bigr\|\sum_{j=0}^{i}Z_{tK+j}\bigr\|$. Then the following hold almost surely for every $0\leq k\leq K$.
@@
    $\sup_{0\leq i\leq k-1}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}\|\leq\alpha\|\nabla F(\tilde{X}_{tK})\|+M_{t,k}.$
@@
    $\sup_{0\leq i\leq k}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK+i}\|\leq\alpha LN_{t}\sup_{i\leq k-1}\|\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}\|.$
@@
    $\mathbb{E}[M_{t,K}^{p}]\leq C(p)(\alpha d)^{\frac{p}{2}}.$
@@
$\sup_{t\leq h\leq s}\left[\mathbb{E}\|\tilde{X}_{hK}-\tilde{X}_{tK}\|^{p}|\tilde{X}_{tK}\right]^{\tfrac{1}{p}}$ $\leq 3\alpha(s-t)\|\nabla F(\tilde{X}_{tK})\|+C_{p}\sqrt{\alpha d(s-t)}.$
@@
$\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}\leq C_{p}L^{p}d^{p}T+C_{p}(\alpha L)^{p-1}\mathbb{E}(\sum_{t=1}^{T}\|\nabla F(\tilde{X}_{tK})\|^{2})^{p}.$ (37)
@@
$\quad+C_{p}L^{p}d^{p}T^{p}+\frac{C_{p}}{\alpha^{p}}.$ (38)
@@
$\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}\leq C_{p}L^{p}d^{p}T(1+(\alpha LT)^{p-1})+\frac{C_{p}L^{p-1}}{\alpha}\left[\mathbb{E}|(F(X_{0})-F(\tilde{X}_{KT}))^{+}|^{p}+1\right].$
@@
    $\mathbb{E}\|B_{tK+i}\|^{2}\leq\frac{CL^{4}\alpha^{5}}{K}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\frac{CL^{4}\alpha^{4}}{K}d.$
@@
    $\mathbb{E}\beta_{tK+i}^{2p}\leq C_{p}\left[L^{2p}\alpha^{3p}K^{p}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2p}+L^{2p}\alpha^{2p}K^{p}d^{p}\right].$
@@
    $\leq\frac{CL^{4}\alpha^{3}}{K}\left[\alpha^{2}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\alpha d\right].$ (39)
@@ E.2 Finishing The Proof
$\quad+L^{10}\alpha^{10}K^{3}d^{5}T+L^{19}\alpha^{19}K^{3}d^{5}T^{5}+L^{14}\alpha^{14}K^{7}d^{7}T+L^{27}\alpha^{27}K^{7}d^{7}T^{7}.$ (40)
@@ Appendix F Underdamped Langevin Dynamics
$A_{h}:=\begin{bmatrix}\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}\\ 0&e^{-\gamma h}\mathbf{I}_{d}\end{bmatrix};\quad G_{h}:=\begin{bmatrix}\tfrac{1}{\gamma}(h-\tfrac{1}{\gamma}(1-e^{-\gamma h}))\mathbf{I}_{d}&0\\ \tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}&0\end{bmatrix};\quad b(X_{t}):=\begin{bmatrix}-\nabla F(U_{t})\\ 0\end{bmatrix}$;
$\Gamma_{h}^{2}:=\begin{bmatrix}\frac{2}{\gamma}\left(h-\frac{2}{\gamma}(1-e^{-\gamma h})+\frac{1}{2\gamma}(1-e^{-2\gamma h})\right)\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-2e^{-\gamma h}+e^{-2\gamma h})\mathbf{I}_{d}\\ \tfrac{1}{\gamma}(1-2e^{-\gamma h}+e^{-2\gamma h})\mathbf{I}_{d}&(1-e^{-2\gamma h})\mathbf{I}_{d}\end{bmatrix}$.
@@
$G_{h}^{\intercal}\Gamma_{h}^{-2}G_{h}\preceq\begin{bmatrix}C\frac{h\exp(2\gamma h)}{\gamma}\mathbf{I}_{d}&0\\ 0&0\end{bmatrix}$.
For all $0\leq j\leq K-1$,
@@ F.1 Bounding the Moments:
Suppose that $\|\cdot\|_{\mathsf{any}}$ is any semi-norm over $\mathbb{R}^{d}$. Then the following hold almost surely for every $0\leq k\leq K$:
@@
    $\quad+\|\sum_{j=0}^{i-1}A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Y_{tK+j}\|_{\mathsf{any}}$
@@
Consider the seminorm given by $\|x\|^{2}_{\Pi}=x^{\intercal}\Pi x$. Let $M_{t,K}:=\sup_{j\leq K-1}\|\sum_{j=0}^{i-1}A_{\tfrac{\alpha(i-1-j)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}\|_{\Pi}$. For any $0\leq i\leq K-1$.
@@
    $\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|^{2}\leq C\alpha^{2}\|\tilde{V}_{tK}\|^{2}+C\alpha^{4}\|\nabla F(\tilde{U}_{tK})\|^{2}+CM_{t,K}^{2}.$ (42)
@@
    $\quad+\frac{N^{2}_{t}\alpha L^{2}}{\gamma}\left(1+\frac{3\alpha L}{2\gamma}\right)\sup_{0\leq i\leq K-1}\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|^{2}.$ (43)
@@
    $\mathbb{E}M_{t,K}^{p}\leq C\exp(\frac{p\gamma\alpha}{2})\gamma^{\tfrac{p}{2}}\alpha^{\tfrac{3p}{2}}(d+\log K)^{\frac{p}{2}}.$
@@
    $\quad+\frac{L^{2}T^{1-\tfrac{1}{p}}\alpha^{2}}{\gamma^{2}}[\mathbb{E}\mathcal{S}_{2p}(V)]^{\tfrac{1}{p}}+\frac{L^{2}T^{1-\tfrac{1}{p}}\alpha^{4}}{\gamma^{2}}[\mathbb{E}\mathcal{S}_{2p}(\nabla F)]^{\tfrac{1}{p}}.$ (44)
@@
    $\quad+C_{p}T^{1-\tfrac{1}{p}}L^{2}\alpha^{2}[\mathbb{E}\mathcal{S}_{2p}(V)]^{\tfrac{1}{p}}+C_{p}T^{1-\tfrac{1}{p}}L^{2}\alpha^{4}[\mathbb{E}\mathcal{S}_{2p}(\nabla F)]^{\tfrac{1}{p}}.$ (45)
@@
    $\quad+\frac{L^{2}T^{1-\tfrac{1}{p}}\alpha^{4}}{\gamma^{2}}[\mathbb{E}\mathcal{S}_{2p}(\nabla F)]^{\tfrac{1}{p}}.$ (46)
@@
    $\quad+C_{p}T^{1-\tfrac{1}{p}}L^{2}\alpha^{2}[\mathbb{E}\mathcal{S}_{2p}(V)]^{\tfrac{1}{p}}+C_{p}T^{1-\tfrac{1}{p}}L^{2}\alpha^{4}[\mathbb{E}\mathcal{S}_{2p}(\nabla F)]^{\tfrac{1}{p}}.$ (47)
@@
$\quad+C_{p}\sqrt{\frac{d+\log K}{L}}.$ (48)
@@
$\quad+C_{p}\gamma\sqrt{\frac{d+\log K}{L}}.$ (49)
@@
    $\mathbb{E}\mathcal{S}_{2p}(\nabla F)\leq C_{p}(L\alpha N)^{2p}\mathbb{E}\mathcal{S}_{2p}(V)+C_{p}TL^{p}\left(d+\log K\right)^{p}+\frac{2^{p}}{N^{p-1}}\mathbb{E}(\mathcal{S}_{2}(\nabla F))^{p}.$ (50)
@@
    $\mathbb{E}\mathcal{S}_{2p}(V)\leq C_{p}(\alpha N)^{2p}\mathbb{E}\mathcal{S}_{2p}(\nabla F)+T\left[\frac{\gamma^{2}}{L}(d+\log K)\right]^{p}+\frac{2^{p}}{N^{p-1}}\mathbb{E}(\mathcal{S}_{2}(V))^{p}.$ (51)
@@
$\quad+C_{p}NTL^{p}\left(d+\log K\right)^{p}.$ (52)
@@
$\mathbb{E}\mathcal{S}_{2p}(\nabla F)\leq C_{p}(L\alpha N)^{2p}\mathbb{E}\mathcal{S}_{2p}(V)+C_{p}TL^{p}\left(d+\log K\right)^{p}+\frac{2^{p}}{N^{p-1}}\mathbb{E}(\mathcal{S}_{2}(\nabla F))^{p}.$ (53)
@@
$\quad+C_{p}NT\frac{\gamma^{2p}}{L^{p}}\left(d+\log K\right)^{p}.$ (54)
@@
$\mathbb{E}\mathcal{S}_{2p}(V)\leq C_{p}(\alpha N)^{2p}\mathbb{E}\mathcal{S}_{2p}(\nabla F)+T\left[\frac{\gamma^{2}}{L}(d+\log K)\right]^{p}+\frac{2^{p}}{N^{p-1}}\mathbb{E}(\mathcal{S}_{2}(V))^{p}.$ (55)
@@
$\quad C_{p}T\left[\tfrac{\gamma^{4p}}{L^{p}}+(\gamma\alpha T)^{p-1}\gamma^{2p}\right](d+\log K)^{p}.$
@@
$\quad+C_{p}T\left[\tfrac{\gamma^{2p}}{L^{p}}+(\gamma\alpha T)^{p-1}\right](d+\log K)^{p}.$